A person starts from rest at the top of a large fr
Answer #1

Q2

Person starts from rest, that is initial speed is zero V_{i}=0

At a point where radius makes an angle \theta with vertical let the speed be V.

conserving energy of person

mgR+\frac{1}{2}mV_i^2=mgR\cos{\theta}+\frac{1}{2}mV^2

V^2=2gR(1-\cos{\theta})

V=\sqrt{2gR(1-\cos{\theta})}

At this point forces on the person are gravitational force mg and Normal reaction N.

net force towards the center is giving the centripetal acceleration mg\cos{\theta}-N=\frac{mV^2}{R}

(a)

If the person is about to lose contact at this point Normal reaction is N=0. That is mg\cos{\theta}=\frac{mV^2}{R}

Hence V^2=gR\cos{\theta}

from earlier equation V^2=2gR(1-\cos{\theta})

Using the above equations \cos{\theta}=\frac{2}{3}

That is the person leaves the surface at an angle \theta=\arccos{\frac{2}{3}}=48.2^{\circ}

(b)

Just after the person leaves contact his speed is V=\sqrt{\frac{2gR}{3}}

From this point his displacement to reach the water is h=\frac{2R}{3}

conserving energy again,

\frac{1}{2}mV_{f}^2=\frac{2mgR}{3}+\frac{1}{2}mV^2

\frac{1}{2}mV_{f}^2=\frac{2mgR}{3}+\frac{mgR}{3}=mgR

V_{f}=\sqrt{2gR}